The table describes hypothetical employment statistics in the United States in a given year. The labor force is the sum of those that are employed plus the unemployed. Labor Full-time workers and UnderemployedDiscouraged force voluntary part-workers 155 million workers workers 140 million 7 mion 5 mon Based on the information, what is the unemployment rate? Specify your answer to one decimal place. unemployment rate: 5.2
Answer #1

Unemployed Laber Farce 100 UnemploymentRate -

Unemployed = Labour Force - Employed Force

Unemployed = 155 - ( 140 + 7) = 155 - 147 = 8 million

Labour force = 155

Unemployment Rate--* 100 155

Unemployment rate = 5.16%

(Note: Here already discouraged worker is excluded so need not to include it in calculation.

Employed = Working + Looking for work)

Please contact if having any query thank you.